OpenStudy (anonymous):

A family from Hong Kong is visiting the United States. Their first stop is a bank, where the following exchange rate is posted: USD to HKD = 7.78. On this day, how many U.S. dollars will the family get for 10 of their Hong Kong dollars? Round to the nearest cent. @ganeshie8

ganeshie8 (ganeshie8):

\(\large \dfrac{1~ USD}{7.78~ HKD} = 1\)

ganeshie8 (ganeshie8):

\(10~ HKD = 10~ HKD \times \large \dfrac{1~ USD}{7.78~ HKD} = \dfrac{10}{7.78} USD \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

answer is 1.29 :)

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

you could also do this by setting up a proportion...